The estimated annual cash flow of an investment project along with associated
probabilities are given below. Determine the expected present worth of this annual cash
flow series at an interest rate of 11% per year.

Year P = 0.5 P = 0.4 P = 0.1
0 ($57,000) ($59,000) ($57,500)
1 $3100 $3000 $3100
2-6 $3250 $3000 $3100
7 $3400 $6000 $3100
Answer #1

i = 11%

Expected PW = 0.5 x PW1 + 0.4 x PW2 + 0.1 x PW3

PW1 = -57000 + 3100/1.11 + 3250/1.112 + .... + 3250/1.116 + 3400/1.117 = - $ 41748.25

PW2 = -59000 + 3000/1.11 + 3000/1.112 + .... + 3000/1.116 + 6000/1.117 = - $ 43418.436

PW3 = -57500 + 3100/1.11 + 3100/1.112 + .... + 3100/1.117 = - $ 42892.19

Expected PW = - $ 42530.72